The un-damped system has angular frequency w = 15 Hz.

The system is then damped with damping constant y = 5 Hz.

At what angular frequency must the system be driven in order to achieve the maximum amplitude?


1
Expert's answer
2021-05-23T16:34:31-0400

The maximum amplitude results when the frequency of the driving force equals the natural frequency of the system


"\omega=\frac{15}{2\pi}=2.4\frac{rad}{s}"